OpenStudy (cherrilyn):

what's the derivative of 3e^x?

OpenStudy (anonymous):

\[3e^x\]

OpenStudy (anonymous):

\[f(x)=e^x\] is its own derivative. imagine that.

OpenStudy (anonymous):

polpak every hear "i'm my own granpa"?

OpenStudy (anonymous):

Recall that for some constant K and some function f of x: \[\frac{d}{dx} K\cdot f(x) = K \cdot \frac{d}{dx} f(x)\] And that \[\frac{d}{dx} e^x = e^x\] If you let K = 3, and \(f(x) = e^x\) you should be able to come up with the solution.